A 45- 45- 90 triangle has a hypotenuse of length 15 units. What is the length of one of the legs?
Found 2 solutions by rfer, RAY100:
Answer by rfer(16322)

take the sin of 45 degrees times 15=10.6 each leg

Answer by RAY100(1637)

Remember that a 45-45-90 triangle has sides proportional to 1-1-sqrt2
.
sqrt2/15 = 1/ x ,,,or x = (15*1)/ sqrt2 = 10.61 units
.
checking, using pythagorous, x^2 + x^2 = 15^2, 2 x^2 = 15^, 10.61 ok